# Dialectal and Low Resource Machine Translation for Aromanian

### v1

@@ Abstract
We present a neural machine translation system that can translate between Romanian, English, and Aromanian (an endangered Eastern Romance language); the first of its kind. BLEU scores range from 17 to 32 depending on the direction and genre of the text. Alongside, we release the biggest known Aromanian-Romanian bilingual corpus, consisting of 79k cleaned sentence pairs. Additional tools such as an agnostic sentence embedder (used for both text mining and automatic evaluation) and a diacritics converter are also presented. We publicly release our findings and models.111https://huggingface.co/aronlp Finally, we describe the deployment of our quantized model at https://arotranslate.com.
@@
Dialectal and Low Resource Machine Translation for Aromanian
Alexandru-Iulius Jerpelea Alina-Ștefania Rădoi Sergiu Nisioi Tudor Vianu National College of Computer Science, Bucharest West University of Timișoara Human Language Technologies Research Center, University of Bucharest alex.jerpelea@gmail.com, alina.radoi@e-uvt.ro, sergiu.nisioi@unibuc.ro
@@ 1 Introduction
Training good machine translation systems in the low-resource setting is a task far from being solved Wang et al. (2021); Haddow et al. (2022). Current advances in large language models and evaluation methodologies are centered on English or are massively multilingual, but do not engage with the particularities of low-resource languages. Of the 7000+ languages spoken in the world, only a small part is covered by current machine translation (MT) systems Costa-jussà et al. (2022).
Aromanian (ISO 639-3 - rup) is an endangered Campbell et al. (2022) Eastern Romance language that currently lacks large-scale corpora and electronic resources that can potentially contribute to the preservation of its cultural heritage. In this study, we are interested in building a dataset suitable for training machine translation systems in the low-resource setting between two closely related languages from the same dialectal continuum: Romanian (ro) and Aromanian (ISO 639-3 - rup). Given the close similarity Caragiu-Marioțeanu (1975) between the two languages, we hope to expand the current pretrained LLM and machine translation models available for standard Romanian to its Eastern Romance sibling, in order to contribute to the preservation of this endangered222https://endangeredlanguages.com/lang/963, Accessed: 2024-09-16 language.
We believe the difficulties for such a task stem from the fact that: (1) there is little availability of monolingual or bilingual texts, (2) the writing system, vocabulary, and grammar have not been standardized in a widely accepted institutional manner, and several types of spelling are currently being used Caragiu-Marioțeanu (1997); Cunia (1997); Nevaci (2008); Ballamaci (2010), and (3) Aromanian has several varieties / dialects that have been influenced by the contact with Greek, Romanian, Turkish, Albanian, and South Slavic languages Caragiu-Marioțeanu (1975). Furthermore, the language has historically been transmitted orally within families or small communities Gica et al. (2009); Maiden (2016), with public usage limited to a few localities in Albania and North Macedonia.
Aiming to fill this gap, our contribution can be summarized as follows:
the creation of a 79k multigenre parallel Aromanian-Romanian sentence-aligned parallel corpus, augmented with with machine translated English sentences
a LaBSE sentence encoder Feng et al. (2020), fine-tuned for Aromanian support
a comparison of different NLLB-200 Costa-jussà et al. (2022) models fine-tuned to translate in any direction between Aromanian, Romanian, and English.
a range of instruction-tuned large language models (LLMs), fine-tuned for translation
a diacritics converter between the multiple orthographic variations of Aromanian
@@ 2 Related Work
Low resource MT has gained considerable traction in the past few years. Initiatives range from commercial ambitions to build ”machine translation systems for the next thousand languages” Bapna et al. (2022), the NLLB family of models that can translate between more than 200 languages Costa-jussà et al. (2022), to approaches that leverage generative language models Alves et al. (2024); Xu et al. (2023). Options for closely related languages employing transliteration or transformer models have been attempted for Belorussian to Russian and Arabic dialects Karakanta et al. (2018); Kumar et al. (2021). To our knowledge, no successful attempt has been made to integrate Aromanian in a translation system.
In terms of electronic resources for Aromanian, they consist mainly of multilingual word-aligned lists (Nisioi, 2014; Cristea and Dinu, 2020; Beniamine et al., 2020; Fourrier and Sagot, 2022; Mititelu et al., 2021) to study language evolution, language contact, historical linguistics, and the relationship with other Romance languages.
The recent work of Petrariu and Nisioi (2024) presents a multilingual parallel corpus of approximately 3k sentences covering Aromanian. Although this is the largest published resource so far, the corpus only covers one genre, namely fairy tales, and the total number of sentence pairs is relatively small, rendering it insufficient to train qualitative MT models, as concluded by Petrariu and Nisioi (2024) when obtaining a maximum of 8 BLEU points.
Our work is the first to build an end-to-end machine translation system for Aromanian language together with the release of a large multi-genre corpus.
@@ 3 Collecting a Multigenre Dataset
In this section, we describe the extraction and alignment method for each parallel text used to train and test our MT system. These texts pertain to different genres such as news articles, literature, dictionaries, religious texts, music lyrics, and essays. They focus on the Romanian-Aromanian language pair.
@@
The Bible has been translated into Aromanian by Dina Cuvata and was published in 2004. The only available version that the we could find is an online scan of the printed edition.333https://dinitrandu.com/wp-content/uploads/2022/04/Bibliea-limba-armaneasca.pdf, Accessed: 2024-09-16 We manually check each page of the scan and perform OCR with Tesseract V5 Smith (2007) on the pages that are intelligible. Aromanian is not supported by the engine, so we set the language parameter to Romanian. We then carefully check each verse and pair it with its match from the Romanian translated Bible. Each verse is divided into sentences in both languages. If the number of sentences per verse matches, we pair the sentences and add them to our corpus. We obtain around k sentence pairs from this source.
Lyrics Translate444https://lyricstranslate.com, Accesed: 2024-09-16 is a platform where songs are translated into multiple languages. We scrape the 500+ songs in both Aromanian and Romanian versions, and pair them verse by verse, obtaining k pairs.
The Multilingual Parallel Corpus of Aromanian (MPC-rup) consists of approximately 3000 pairs of sentences from Aromanian fairy tales and short prose Petrariu and Nisioi (2024).
A list of parallel short phrases and a set of anonymously authored long sentences and phrases dictionary, consisting of k pairs, which we have received as a donation.
The Avdhela project,555https://www.proiectavdela.ro Accessed: 2024-09-16 is a digital Aromanian library consisting of a collection of parallel poetry and prose. We extract 4000 pairs.
Radio Romania International is a Romanian public news radio station that includes editorials in many languages, among them Aromanian. The majority of radio transmission are published in text form on the official website666https://www.rri.ro/ro_ar, Accessed: 2024-09-16 and Aromanian articles are usually translations of matching Romanian articles.777A fact confirmed by several authors working for the radio station. However, there are no linkbacks to these pages, which makes it difficult to trace back the corresponding articles. Thus, we scrape all articles in both languages and match them by the images they contain, as we believe this to be an efficient proxy. In cases where multiple matches are possible, we deploy our sentence alignment tool (described in section 4) between the set of titles in Aromanian and the set of titles in Romanian, pairing the titles with the greatest semantic similarity, and retaining the possibility of unmatched titles. This method yields k sentence pairs.
The adventures of Tom Sawyer Twain (2007) was translated into Aromanian from Romanian and donated to us (as it is not yet published). We align the sentences with the same alignment tool described in section 4 and yield a total of k sentences.
A century of Aromanian poetry Cândroveanu and Iorgoveanu (1985) is a collection of Aromanian poetry translated into Romanian. We use noisy scanned version and the Open CV Bradski (2000) computer vision tool to identify text bounding boxes (Figure 1), and, considering their location within the image, establish if the text chunk is in Aromanian or Romanian. Next, for each page, we align the resulting sentences using the alignment tool described in section 4, which adds k pairs to the corpus.
@@
The Divine Comedy Dante (2017) was translated by Dina Cuvata in Aromanian888https://dinitrandu.com/wp-content/uploads/2018/08/Dante-Dina-Cuvata.pdf, Accessed: 2024-09-16, and by aligning it with the Romanian version we obtain k pairs.
An anonymously donated collection of modern Aromanian poetry, which cannot be found online and is already aligned, contains k pairs.
Tao Te Ching has been partially translated999https://dinitrandu.com/wp-content/uploads/2022/06/Cartea-a-Calillei.pdf, Accessed: 2024-09-16. By aligning it with its Romanian counterpart, we obtain 260 sentences.
The translated Aromanian version of ”The Little Prince” is aligned similarly to the Romanian version with sentence pairs.
Writings is a collection of news articles consisting of news articles, personal stories, and essays, which are not available online. These were donated to us by Kira Mantsu, a prominent Aromanian writer. The documents are originally in Aromanian and translated into Romanian by the author herself; the alignment process results in a total of k new pairs.
3.1 Dataset Split
Considering all the sources mentioned above, the dataset we use to build our models consists of k sentence pairs. A table with sample sentences from each source can be viewed in Appendix A.
In preparation for the model training process, we split our dataset into train, dev and test splits. Sentences from ”The Little Prince” and Writings are only included in the dev and test sets, but are not used for training. The reason behind this choice is that we want to evaluate the model’s performance on novel data, and we believe that the literary genre is a difficult evaluation task.
For our train set, we extract in a stratified manner random pairs from each of the above sources. The remaining pairs from those sources make up our dev set. We also add the sentences from ”The Little Prince” to our dev set to observe model evaluation scores during training. The test set comprises only Writings documents to enforce evaluation on out-of-domain data; thus, the texts do not have a common authorial origin with the data on which our models are trained. The total number of sentence pairs per source per split can be visualized in Table 1.
@@ 3.1 Dataset Split
| Split | Bible | Lyrics | Small corpus | Phrases | Avdhela | Radio | Poetry | Tom Sawyer |
@@
| train | ||||||
| dev | ||||||
| test |
@@ 3.2 Synthetic Data
Because our aim is to also perform translations between Aromanian and English, we add synthetic English counterparts to all of the sentence pairs by translating Romanian sentences with the Google Translate API.
@@
To evaluate the quality of automatic English translations, we employ the Cometkiwi-da model Rei et al. (2022) for reference-free evaluations, therefore English translations are evaluated only with regard to the Romanian original sentences. Throughout the corpus, we obtain a system-score of (scores range from to ). The resulted score is an indicator that Romanian-to-English machine translations are of mediocre quality and that automatic translations to and from English are error-prone, but still meaningful.
3.3 Dataset Size
@@
It is important to note that there are numerous orthographic standards for Aromanian. The main spelling types in our corpus are DIARO (named after the Aromanian-Romanian Dictionary by Caragiu-Marioțeanu (1997), and Cunia, named after Tiberius Cunia, the author of the Aromanian Dictionary, Cunia (2010). We also sparsely found Greek orthography, which we standardized using the Latin script.
The Cunia and DIARO spellings are easily convertible to each other using simple regular expressions, with the exception of the closed central and mid-central vowels. Both of these sounds are represented by the letter ¡ã¿ in the Cunia standard. However, the DIARO standard follows the Romanian standardization Petrariu and Nisioi (2024) using two different symbols ¡î¿ and ¡â¿ for closed central and ¡ă¿ for mid central.
@@
We chose to convert our entire corpus to the Cunia writing system, the main reason for this choice being the slightly lower fertility rate of the BPE tokenizers versus DIARO orthography that reaches tokens per word. We observed lower fertility scores across various models including both NLLB-200 and pre-trained large language models. Since the pre-trained models are multilingual, it is likely that the tokenizers have seen byte-pairs similar to Cunia orthography in words from Albanian, South Slavic, and Greek Latin transliterated versions.
Combating the potential loss of details due to the merging of the mid-central and closed central vowels, we train an n-gram-based model to convert from Cunia to DIARO.
@@
Table 2 presents several statistics regarding our Cunia-converted datasets. Original Aromanian texts have the greatest lexical richness and a tendency to create shorter phrases. However, automatically aligned human translations into Romanian and the machine translated texts in English have both lower lexical diversity (as estimated by type-token ratio) and larger sentence lengths. In Appendix B, we also present these statistics for each text genre separately.

### v2

# Dialectal and Low-Resource Machine Translation for Aromanian

@@ Abstract
This paper presents the process of building a neural machine translation system with support for English, Romanian, and Aromanian - an endangered Eastern Romance language. The primary contribution of this research is twofold: (1) the creation of the most extensive Aromanian-Romanian parallel corpus to date, consisting of 79,000 sentence pairs, and (2) the development and comparative analysis of several machine translation models optimized for Aromanian. To accomplish this, we introduce a suite of auxiliary tools, including a language-agnostic sentence embedding model for text mining and automated evaluation, complemented by a diacritics conversion system for different writing standards. This research brings contributions to both computational linguistics and language preservation efforts by establishing essential resources for a historically under-resourced language. All datasets, trained models, and associated tools are public:111https://huggingface.co/aronlp https://arotranslate.com
@@
Dialectal and Low-Resource Machine Translation for Aromanian
Alexandru-Iulius Jerpelea Alina Rădoi Sergiu Nisioi††thanks: Corresponding author. Tudor Vianu National College of Computer Science, Bucharest West University of Timișoara Human Language Technologies Research Center, University of Bucharest alex.jerpelea@gmail.com, alina.radoi@e-uvt.ro, sergiu.nisioi@unibuc.ro
@@ 1 Introduction
Training good machine translation (MT) systems in a low-resource setting is a task far from being solved Wang et al. (2021); Haddow et al. (2022). Current advances in Large Language Models (LLMs) and evaluation methodologies are centred on English or are massively multilingual, but do not engage with the particularities of low-resource languages. Of the 7000+ languages spoken in the world, only a small part is covered by current MT systems Costa-jussà et al. (2022).
Aromanian (ISO 639-3 - rup) is an endangered Eastern Romance language Moseley and Nicolas (2010), which currently lacks large-scale corpora and electronic resources that can potentially contribute to the preservation of its cultural heritage. In this study, we build a dataset suitable for training MT systems in the low-resource setting between two related languages from the same dialect continuum: Romanian (ron) and Aromanian (ISO 639-3 - rup). Given the close similarity Caragiu-Marioțeanu (1975) between the two languages, we expand the existing pre-trained LLMs and machine translation models available for standard Romanian to its Eastern Romance sibling, in order to contribute to the preservation of this endangered language.222https://endangeredlanguages.com/lang/963, Accessed: 2024-09-16
The difficulties for such a task stem from the fact that: (1) there is little availability of monolingual or bilingual texts, (2) the writing system, vocabulary, and grammar have not been standardized in a widely accepted institutional manner, several types of spelling are currently being used Caragiu-Marioțeanu (1997); Cunia (1997); Nevaci (2008), and (3) Aromanian has several varieties / dialects that have been influenced by contact with Greek, Romanian, Turkish, Albanian, and South Slavic languages Caragiu-Marioțeanu (1975); Pascaru and Kahl (2017). Furthermore, the language has historically been transmitted orally within families or small communities Gica et al. (2009); Maiden (2016), with public usage limited to a few towns in Albania and North Macedonia.
Our contribution is:
the creation of a 79k multigenre Aromanian-Romanian sentence-aligned parallel corpus, augmented with machine-translated English sentences;
a language-agnostic BERT Sentence Encoder (LaBSE) sentence encoder Feng et al. (2022), fine-tuned for Aromanian support;
a comparison of different No Language Left Behind (NLLB) Costa-jussà et al. (2022) models fine-tuned to translate in any direction between Aromanian, Romanian, and English;
a range of instruction-tuned large language models, fine-tuned for Aromanian translation;
a diacritics converter between two major orthographic variants of Aromanian.
@@ 2 Related Work
Low-resource machine translation has gained considerable traction in the past few years. Initiatives range from commercial ambitions to build “machine translation systems for the next thousand languages” Bapna et al. (2022), the NLLB family of models that can translate between more than 200 languages Costa-jussà et al. (2022), to approaches focusing on severely under-resourced languages Parida et al. (2021); Sánchez-Martínez et al. (2024).
Varieties and closely related languages also receive special attention, from early work on dialectal and language varieties translations exploring statistical machine translation and rule-based systems Zhang (1998); Scannell (2006); Otte and Tyers (2011); Hassani (2017) to more recent WMT Shared Tasks Akhbardeh et al. (2021) covering translations from the same language families, such as Dravidian, Manding, and Romance languages. Although rule-based methods are still a strong baseline for certain language pairs Sánchez-Martínez et al. (2024), neural approaches provide the state of the art in multiple tasks from Portuguese Costa-jussà et al. (2018), Serbo-Croatian Popović et al. (2020), Belorussian-Russian, to Arabic dialects Karakanta et al. (2018); Kumar et al. (2021).
Aromanian remains an unexplored language in MT research. The electronic resources available for Aromanian consist primarily of multilingual word-aligned lists (Nisioi, 2014; Cristea and Dinu, 2020; Beniamine et al., 2020; Mititelu et al., 2021; Fourrier and Sagot, 2022), which are used to study its evolution, history, contact relationship with other Romance languages. The recent work of Petrariu and Nisioi (2024) presents a multilingual parallel corpus of approximately 3k sentences covering Aromanian. While it is the largest published resource to date, this corpus only covers one genre, namely fairy tales, and the total number of sentence pairs renders it insufficient for training qualitative MT models, as concluded by Petrariu and Nisioi (2024).
To the best of our knowledge, no successful attempts have been made to integrate Aromanian into a translation system, and our work lays the first building blocks in this direction.
@@ 3 Collecting a Multigenre Dataset
The original texts collected in our dataset focus on the Romanian-Aromanian language pair and pertain to different genres such as news articles, literature, dictionaries, religious texts, music lyrics, and essays. Throughout the data collection process, we use a sentence embedder compatible with Aromanian to align sentences across languages. We fine-tune a language-agnostic BERT Sentence Encoder Feng et al. (2022) using a similar approach to Dale (2022) (see Section 4).
@@
The Bible has been translated by Dina Cuvata and was published in 2004. We use an online scan of the printed edition published by the Aromanian Library hosted by Dini Trandu.333https://dinitrandu.com/wp-content/uploads/2022/04/Bibliea-limba-armaneasca.pdf, Accessed: 2024-09-16 Each page of the scan is digitized using Tesseract V5 Smith (2007). The engine does not support Aromanian, so the language parameter is set to Romanian. Although the OCR quality is mediocre, it is high enough for us to post-process the results. We then carefully check each verse and pair it with its match from the Romanian translated Bible. Each verse is sentence-split using regular expressions in both languages. If the number of sentences per verse matches, we pair the sentences and add them to our corpus. We obtain around k sentence pairs from this source.
@@
The Divine Comedy has been translated by Dina Cuvata from Romanian.444https://dinitrandu.com/wp-content/uploads/2018/08/Dante-Dina-Cuvata.pdf, Accessed: 2024-09-16 By aligning it with the Romanian version, using multilingual sentence embeddings described in Section 4, we obtain k pairs.
Tao Te Ching has been partially translated from English by Mihali Prefti and was published by the Aromanian Library hosted by Dini Trandu.555https://dinitrandu.com/wp-content/uploads/2022/06/Cartea-a-Calillei.pdf, Accessed: 2024-09-16 Using automatic sentence alignment based on LaBSE (see Section 4), we obtain 260 sentences.
Lyrics Translate666https://lyricstranslate.com, Accessed: 2024-09-16 is a platform where songs are translated into multiple languages. We scrape the 500+ songs in both Aromanian and Romanian and pair them verse by verse, obtaining k pairs.
The Multilingual Parallel Corpus of Aromanian (MPC-rup) consists of approximately k pairs of sentences from Aromanian fairy tales and short prose texts Petrariu and Nisioi (2024).
A list of parallel common phrases and idioms, consisting of k pairs, provided for this project by members of the Aromanian community.
The Avdhela Project777https://www.proiectavdela.ro Accessed: 2024-09-16 is a digital Aromanian library consisting of a collection of parallel poetry and prose texts. We use only the poems and pair them verse by verse, thus obtaining another k pairs for our corpus.
Radio Romania International is a Romanian public news radio station that includes editorials in many languages, among them Aromanian. Most radio broadcasts are published in text form on the official website888https://www.rri.ro/ro_ar, Accessed: 2024-09-16 and Aromanian articles are usually translations of matching Romanian articles.999A fact confirmed by several authors working for the radio station. However, there are no backlinks to these pages, making it difficult to trace the corresponding articles. Thus, articles in both languages are scraped and matched by the images they contain. In cases where multiple matches are possible, we use our sentence alignment tool (described in Section 4) to align the titles in Aromanian with those in Romanian, pairing the titles with the greatest semantic similarity while allowing for unmatched titles. Then, for the sentence alignment of each pair of articles, we once again deploy the aforementioned alignment tool. This method yields k sentence pairs.
The Adventures of Tom Sawyer has been translated from Romanian and the digital version was donated for this project. A total of k sentences are obtained after automatic alignment.
A century of Aromanian poetry Cândroveanu and Iorgoveanu (1985) is a collection of Aromanian poetry translated into Romanian. We use a scanned digitized version and apply Open CV Bradski (2000) to identify text bounding boxes (Figure 1). We then apply OCR and the same post-processing steps used for the Bible. Finally, we align the resulting sentences for each page using the alignment tool described in Section 4, adding approximately k pairs to the corpus.
A collection of modern Aromanian poetry written by George Vrana was donated for this project with the author’s agreement. Aligning the text verse by verse resulted in k pairs.
@@
| Split | Bible | Lyrics | MPC-rup | Phrases | Avdhela | Radio | Poetry | Tom Sawyer |
@@
| train | - | - | ||||
| dev | - | |||||
| test | - | - | - | - |
The Little Prince has been translated into Aromanian by Maria Bara and Thede Kahl. We extract sentence pairs using automatic alignment.
Writings is a collection of texts consisting of news articles, personal stories, and essays. These are not available online and were donated for this project by Kira Mantsu, a prominent Aromanian writer. The documents are originally in Aromanian and translated into Romanian by the author herself; the automatic alignment results in a total of k new pairs.
3.1 Dataset Split
Considering all the sources mentioned above, the dataset we use to build our models consists of k sentence pairs. A table with sample sentences from each source is available in Appendix A.
In preparation for the model training process, we split our dataset into train, dev and test splits. Sentences from “The Little Prince” and “Writings” are only included in the dev and test sets, but are not used for training. The reason behind this choice is that we want to evaluate the model’s performance on novel data, and we believe that the literary genre is a difficult evaluation task.
For the train set, we extract in a stratified manner random pairs from each of the above sources. The remaining pairs from these sources make up the dev set. We also add the sentences from “The Little Prince” to the dev set to provide a difficult evaluation set of literary translations. The test set comprises only of the mixed-genre “Writings” documents to enforce evaluation on out-of-domain data; thus, the texts do not have a common authorial origin with the data on which our models are trained. Table 1 contains the exact number of sentence pairs extracted from each source.
@@ 3.2 Synthetic Data
To perform translations between Aromanian and English, we add synthetic English counterparts to all sentence pairs by translating Romanian sentences with the Google Translate API.
@@
To evaluate the quality of automatic English translations, we employ cometkiwi-da model Rei et al. (2020) for reference-free evaluations; therefore, English translations are evaluated only with regard to the original Romanian sentences. Throughout the corpus, we obtain a system-score of (scores range from to ). The resulted score is an indicator that Romanian-to-English machine translations are of mediocre quality and that automatic translations to and from English are error-prone but still meaningful.
3.3 Orthography
@@
It is important to note that there are numerous orthographic standards for Aromanian. The main spelling types in our corpus are DIARO (named after the Aromanian-Romanian Dictionary by Caragiu-Marioțeanu (1997), and Cunia, named after the author of the Cunia (2010) Dictionary. We also sparsely found Greek orthography, which we standardized using the Latin script.
The Cunia and DIARO spellings are easily convertible to each other using simple regular expressions, with the exception of the close central and mid central vowels. Both of these sounds are represented by the grapheme ¡ã¿ (U+00E3) within the Cunia standard. However, the DIARO standard follows the Romanian standardization Petrariu and Nisioi (2024).101010The Romanian standardization remains inconsistently adopted, particularly in online texts. The primary variation is the alternation between ¡â¿ (used mid-word) and ¡î¿ (used at the beginning and end of words) to represent the close central vowel. The introduction of the grapheme ¡â¿ by the Romanian Academy in 1993 as an anticommunist measure has been widely criticized by linguists for its lack of scientific and etymological justification Dumistrăcel (1993). Consequently, certain Romanian publications and publishers allow their authors to choose their preferred standard rather than enforcing this rule. using two different symbols ¡î¿ (U+00EE) and ¡â¿ (U+00E2) for close central and ¡ă¿ (U+0103) for mid central.
@@
We release the corpus in both writing standards, but for training purposes the Cunia writing system is used. The main reason for this choice is the slightly lower fertility rate of the tokenizers versus the DIARO orthography that reaches tokens per word. We observed lower fertility scores across various models, including both NLLB-200 and pre-trained large language models. Since the pre-trained models are multilingual, it is likely that the tokenizers have seen byte-pairs similar to the Cunia orthography in its pretraining data, i.e., words from Albanian and romanised transliterations from South Slavic and Greek languages.
@@
Combating the potential loss of details due to the merging of the mid central and close central vowels, we train an n-gram-based model to convert from Cunia to DIARO. Table 2 presents several statistics regarding our Cunia-converted datasets. Original Aromanian texts have the greatest lexical richness and shorter phrases. However, the Romanian human translations and the machine- translated texts in English have both lower lexical diversity (as estimated by the type-token ratio) and larger sentence lengths. In Appendix B, we also present these statistics for each text genre separately.